OpenStudy (anonymous):

solve: x^2 - 6x + 5 = 0

OpenStudy (anonymous):

x=5, or x=1.

OpenStudy (anonymous):

Factoring: (x - 5)(x - 1) = 0 Zero-product property: x - 5 = 0 and x - 1 = 0 x = {5,1}